OpenStudy (studygurl14):

calc @michele_laino @pooja195 @directrix

OpenStudy (studygurl14):

@mathmale

OpenStudy (studygurl14):

@mathmale @jigglypuff314

OpenStudy (studygurl14):

@mathmale @Directrix @IrishBoy123

jigglypuff314 (jigglypuff314):

First step is to either draw a picture of it to help you visualize what you are looking for or if you just want numbers, find the points where the two given equations intersect.

OpenStudy (studygurl14):

ok

OpenStudy (studygurl14):

after I do that, what next?

jigglypuff314 (jigglypuff314):

what points did you get? :)

OpenStudy (studygurl14):

(1,2) and (-0.75, -1.5)

jigglypuff314 (jigglypuff314):

great! so \[area = \int\limits_{-1.5}^{2}(Upper~equation - Lower~equation)~dy\]

jigglypuff314 (jigglypuff314):

for the upper and lower equations (since we are trying to integrate with respect to y) we should plug in the equations as x = y^2 - 3 and x = 0.5 y

OpenStudy (studygurl14):

so do I just use \(\large\sqrt{x+3}\) rather than \(\large\pm\sqrt{x+3}\)

OpenStudy (studygurl14):

Oh! I see

jigglypuff314 (jigglypuff314):

yeah, so we are plugging in y^2 - 3 and 0.5 y :)

OpenStudy (studygurl14):

what do i use as the limits

OpenStudy (studygurl14):

the x values or the y

jigglypuff314 (jigglypuff314):

use the y values we found of the intersection points so -1.5 and 2 :)

OpenStudy (studygurl14):

ok thank you

OpenStudy (studygurl14):

I got 89/48